The work done on a reversed heat engine is 100kg whilst the heat transfer to the engine
from the low temperature is 300kJ. Determine the heat transfer to the high temperature reservoir and the COP as a refrigerator and as a heat pump.
1
Expert's answer
2020-08-26T11:08:44-0400

Given:

"W=100\:\rm kJ"

"Q_L=300\:\rm kJ"

The the heat transfer to the high temperature reservoir 


"Q_H=Q_L+W=300\:\rm kJ+100\: kJ=400\: kJ"

The COP of a refrigerator

"\rm COP_r=\frac{Q_L}{W}=\frac{300\:\rm kJ}{100\:\rm kJ}=3"

The COP of a heat pump

"\rm COP_{hp}=\frac{Q_H}{W}=\frac{400\:\rm kJ}{100\:\rm kJ}=4"